OpenStudy (anonymous):

What value of h makes the statement true? (17 + h) + (–9) = 17 + (5 + (–9))

OpenStudy (anonymous):

9 because 17 + 9 = 26 and 26 + (-9) is 17

OpenStudy (anonymous):

Associative property of addition \[(a+b) + c = a+(b+c)\] \[(17+h) + (-9) = 17+(5+(-9))\] so h will be?

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

correct

OpenStudy (asett):

Move all the numbers to one side of the equation, leaving the variable on the other: \[h=9-17+17+(5+(-9))\] Solve for h: \[h=5\]

OpenStudy (anonymous):

Ok, thank you all.

OpenStudy (anonymous):

anytime:)

OpenStudy (asett):

No problem!